OpenStudy (refrac532):

>> IMPORTANT << This is homework that is due tomorrow, need help verifying my answers! For the following questions, the parent function is y = x(x-1)(x+4). Describe the transformation that transforms the parent graph into the graph of each function. 1. y=3x(x-1)(x+4) 2. y=x(x/2-1)(x/2+4) (For this question, x/2 is a fraction, and the -1 and +4 come after the fraction.)

OpenStudy (refrac532):

For #1, I got "stretched vertically by a factor of 3". For #2, I got "stretched horizontally by a factor of 2". Please verify these answers

OpenStudy (mathmate):

#1 is correct. For #2, rewrite f(x)=x(x-1)(x+4) y=2(x/2)(x/2-1)(x/2+4) =2f(x/2) which tells us that there is a horizontal stretch factor of 2, AND a vertical stretch factor of 2. Check using x=5, y(5)=195/4 2f(5/2)=195/4

OpenStudy (refrac532):

Alright. Thank you!